# Dissipating quartets of excitations in a superconducting circuit

### v1

@@ II.2 Modes dispersion against bias fluxes
In order to characterize the system, we record the modes resonance frequencies against DC magnetic fluxes biasing the ATS loops in absence of pump ($\epsilon(t)=0$). We first monitor the readout mode which is directly probed in single-tone spectroscopy through the charge line. Due to the weak participation of the ATS, the mode frequency dispersion and non-linearity are weak enough that it remains visible at most flux biases, yielding the frequency map represented in Fig. 2b. We note that in the regime $E_{J}{\varphi^{ZPF}_{a,b}}^{2}\simeq\hbar\omega_{a,b}$ we operate in (see Table 1), this map qualitatively differs from the egg-box pattern one would naively expect if the SQUID contribution to the system Hamiltonian (2) were a small perturbation. Nevertheless, the perturbative picture is valid nearby the ATS operating points where the SQUID inductive energy vanishes (see Eq. (1) and following comment). We identify three of these points as saddle points of the frequency map, allowing us to relate the direct currents applied through the flux lines to the actual fluxes threading the ATS (Fig. 2b uses calibrated flux coordinates  [^4] The fluxes $\Phi_{1}$ and $\Phi_{2}$ are known up to an integer number of flux quanta and depend on applied currents through an affine transformation. The map is centered at non-zero applied currents, indicating significant flux offsets., and two non-equivalent saddle points are identified by crossing arrows). Next, we record the memory and buffer resonance frequencies—respectively in two-tone spectroscopy leveraging the memory-readout dispersive coupling and in single-tone spectroscopy—around the saddle points. Fig. 2c features each mode resonance frequency (empty circles) when sweeping applied fluxes through the saddle points of the frequency map, along directions of maximum and minimum curvatures (indicated by double-headed arrows with matching colors in Fig. 2b). Strikingly, the modes have near identical resonance frequencies at both saddle points (markers in bright colors corresponding to the saddle point $\mathrm{SP}_{1}$ are aligned with markers in dim colors corresponding to the saddle point $\mathrm{SP}_{2}$), indicating that the Josephson junctions forming the SQUID have a relative energy difference smaller than $0.01~{}\%$ [47]. We attribute this exceptionally low circuit asymmetry to chance in fabrication and junction aging behavior. Indeed, the asymmetry was found to be two orders of magnitude higher in similar circuits that we have characterized, and an order of magnitude higher during a previous cooldown of the same circuit.
@@ II.3 Stray inductances and memory non-linearity
Unexpectedly, when driving the memory on resonance and subsequently probing the readout resonator, we detect a reflected signal that oscillates with the duration and amplitude of the memory drive pulse [47]. Measuring such Rabi oscillations unambiguously reveals a significant non-linearity of the memory, whose transitions can be addressed individually. Driving the memory with sequences of charge drives (to excite single-photon transitions $|n\rangle_{a}\leftrightarrow|n+1\rangle_{a}$) and/or flux drives (to excite three-photon transitions $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ leveraging the wave-mixing properties of the ATS), we reconstruct the memory mode energy level structure at the saddle point (see Fig. 2d). The anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$. This pattern is not the one expected for non-linearity induced by the small asymmetry of the SQUID (anharmonicity predicted to be orders of magnitude weaker and with opposite sign at both saddle points [47]) and from the non-linearity of the circuit superinductors (negative anharmonicity near constant with $n$ [47]). On the other hand, the pattern is well captured by a simple model accounting for a small stray inductance in series with the junction on each arm of the SQUID (inductive energy $\epsilon_{L}\gg E_{J}$, see Fig. 2a). Indeed, series inductances renormalize the SQUID potential as $\tilde{V}_{\mathrm{SQUID}}(\varphi)=-E_{J}f(\varphi-2\pi\frac{\Phi_{1}}{\Phi_{0}})-E_{J}f(\varphi+2\pi\frac{\Phi_{2}}{\Phi_{0}})$ where $f(\varphi)=\cos(\varphi)-\frac{E_{J}}{4\epsilon_{L}}\cos(2\varphi)+\mathcal{O}((\frac{E_{J}}{\epsilon_{L}})^{2})$ [53, 54]. The $\cos(2\varphi)$ contribution corresponds to the coherent tunneling of pairs of Cooper pairs. It is maximal around the ATS saddle points and yields a positive anharmonicity that decreases rapidly with the level index $n$ due to the effective doubling of the modes phase-fluctuations [53, 47].
Overall, we fit all data presented in Fig. 2b-d with the energy level structure of the effective circuit pictured in Fig. 2a. The fitted parameters are close to design values (see Table 1). In particular, we note that the memory energy spectrum is well reproduced with a $\cos(2\varphi)$ contribution to the SQUID potential at the level of 0.2 % of the standard $\cos(\varphi)$ term. This is an order of magnitude smaller than the values reported in a recent analysis of transmon energy level structure [55] and can be entirely attributed to the mechanism described above involving series inductances of $\Phi_{0}^{2}/(4\pi^{2}\epsilon_{L})=85$ pH, compatible with the estimated geometric inductance of the circuit. Thus, we do not find signatures of intrinsic distortion of the junctions potential as in [55] which may be explained by a thicker oxide barrier in our junctions [47]. Geometric inductances could be significantly reduced in a future iteration of the experiment by placing the ATS shunt inductor outside of the SQUID loop and scaling down the latter from its current $\sim 50\times 50~{}\mathrm{\mu m}^{2}$ dimensions.
@@ III.1 Preparing a Fock state and measuring its occupation
In detail, the states are prepared with pulse sequences combining $\pi$-pulses applied on the single-photon transitions $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|2\rangle_{a}$ (direct charge drive) and/or on the three-photon transitions $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|4\rangle_{a}$ (flux drive of the ATS). The latter are favored over sequences of pulses on single-photon transitions as the detuning between neighboring three-photon transitions is larger, allowing for stronger drives. Yet, all pulses have a duration on par with the coherence time of the transitions, so that preparation is imperfect. Nevertheless, only energy states $|k\rangle_{a}$ with $k\leqslant n$ are occupied when preparing $|n\rangle_{a}$. This allows us to subsequently measure the occupation of $|n\rangle_{a}$ up to an unknown proportionality factor, as we now detail.
@@
The principle of this measurement is depicted in Fig. 3a. While the memory mode vacuum state $|0\rangle_{a}$ exerts a dispersive pull onto the readout mode which is significantly different from the pull exerted by excited states, the differential pull exerted by one excited state and the next tails off rapidly. This makes excited states hard to distinguish directly. We circumvent this shortcoming by applying or not a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long pulse. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the vacuum state is delayed compared to the decay from $|n\rangle_{a}$, we detect a time-shifted signal in the reflected field $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse with respect to the field $s_{n}^{OFF}(t)$ detected in absence of the pulse [47]. We integrate the detected fields with a time-envelope $\mathcal{E}(t)$ chosen to maximize the signal-to-noise ratio of the differential signal $S_{n}=\int_{t}\Re\Big{(}\big{(}s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big{)}\mathcal{E}(t)\Big{)}\mathrm{d}t$ between these two situations [47]. Crucially, the $\pi$-pulse only modifies the memory state when $|n\rangle_{a}$ is occupied so that, even though it does not perfectly transfer population from $|n\rangle_{a}$ to $|n+3\rangle_{a}$, the integrated differential signal $S_{n}$ is proportional to the occupation of state $|n\rangle_{a}$ at the instant when the pulse is delivered. Note that we chose to apply the on/off pulse on a three-photon transition rather than a single-photon transition for similar reasons as for state preparation, and to maximize the differential signal.
@@
Before applying these preparation and measurement methods to characterize the memory population dynamics in presence of a pump, we report that some of the memory transitions—driven either during preparation or measurement, involving a single photon or three—were frequently and transiently observed to shift or split [47]. We attribute this effect to hybridization of the memory to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [57, 58]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, the probability that at least one of the memory transitions be affected by the TLS bath was high enough that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the memory transitions away from any resonant coupling with TLS’s [47]. The frequency of this drive is chosen to minimize decoherence induced on the memory, but still leads to degradation of its energy relaxation and dephasing times. The datasets presented in the following section were acquired in presence of such a drive.
@@ III.2 Characterization of the four-photon dissipation channel
In order to quantitatively relate the largest fitted values of $g_{4}$ to four-photon dissipation rates, one needs to go beyond the simple model reduction allowed by adiabatic elimination of the buffer. To this end, we derive from the two-mode Lindblad master equation an evolution matrix encoding the dynamics of the expectation values of a set of operators, among which $|4\rangle_{a}\langle 4|_{a}$ [47]. Through spectral analysis, we then identify the matrix eigenvalue corresponding to the effective decay $\Gamma_{4}$ of the population in state $|4\rangle_{a}$, and report its value in Fig. 4 (filled red circles). Following an initial quadratic increase of $\Gamma_{4}$ against pump amplitude—also predicted by the adiabatic elimination of the buffer—we observe a saturation of $\Gamma_{4}$ near its theoretical maximum value $\kappa_{b}/2$ set by the buffer damping rate [23]. In this regime, engineered four-photon dissipation dominates the dynamics, inducing decay of population from state $|4\rangle_{a}$ at a rate that exceeds that induced by native single-photon dissipation by a factor 10. As a sanity check, we also record, under the same conditions, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ and report the corresponding decay rates $\Gamma_{n}$ in Fig. 4. While these should nominally not be affected by the pump (see Fig. 1b), we observe a small but clear increase with pump amplitude. Such behavior was reported in similar parametrically pumped systems [19, 21], and are intensively investigated [59, 60, 54, 61]. In any case, the spurious increase of $\Gamma_{n}$ for $n<4$ remains an order of magnitude lower than the recorded increase of $\Gamma_{4}$.
IV Conclusion and outlook

### v2

@@ II.2 Modes dispersion against bias fluxes
In order to characterize the system, we record the modes resonance frequencies against DC magnetic fluxes biasing the ATS loops in absence of pump ($\epsilon(t)=0$). We first probe the readout mode in single-tone spectroscopy through the charge line, yielding the frequency map represented in Fig. 2b. We identify the ATS working point corresponding to the DC biases introduced in Eq. (1) as a saddle point of the frequency map, along with analog points located at half-integer numbers of flux quanta away. This allows us to relate the direct currents applied through the flux lines to the actual fluxes threading the ATS (Fig. 2b uses calibrated flux coordinates  [^3] The fluxes $\Phi_{1}$ and $\Phi_{2}$ are known up to an integer number of flux quanta and depend on applied currents through an affine transformation. The map is centered at non-zero applied currents, indicating significant flux offsets., and two non-equivalent saddle points are identified by crossing arrows). Next, we record the memory and buffer resonance frequencies—respectively in two-tone spectroscopy leveraging the memory-readout dispersive coupling and in single-tone spectroscopy—around the saddle points. Fig. 2c features each mode resonance frequency (empty circles) when sweeping applied fluxes through the saddle points of the frequency map, along directions of maximum and minimum curvatures (indicated by double-headed arrows with matching colors in Fig. 2b). Strikingly, the modes have near identical resonance frequencies at both saddle points (markers in bright colors corresponding to the saddle point $\mathrm{SP}_{1}$ are aligned with markers in dim colors corresponding to the saddle point $\mathrm{SP}_{2}$), indicating that the Josephson junctions forming the SQUID have a relative energy difference smaller than $0.01~{}\%$ [40]. We attribute this exceptionally low circuit asymmetry to chance in fabrication and junction aging behavior. Indeed, the asymmetry was found to be two orders of magnitude higher in similar circuits that we have characterized, and an order of magnitude higher during a previous cooldown of the same circuit.
@@ II.3 Stray inductances and memory non-linearity
Unexpectedly, when driving the memory on resonance and subsequently probing the readout resonator, we detect a reflected signal that oscillates with the duration and amplitude of the memory drive pulse [40]. Measuring such Rabi oscillations unambiguously reveals a significant non-linearity of the memory, whose transitions can be addressed individually. Driving the memory with sequences of charge drives (to excite single-photon transitions $|n\rangle_{a}\leftrightarrow|n+1\rangle_{a}$) and/or flux drives (to excite three-photon transitions $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ leveraging the wave-mixing properties of the ATS), we reconstruct the memory mode energy level structure at the saddle point (see Fig. 2d). The anharmonicity $\omega_{a}^{n+1,n+2}-\omega_{a}^{n,n+1}$ is positive for all $n$, lies in the 1—10 MHz range and decreases sharply with $n$. As discussed in Sec. IV, we unambiguously relate this exotic pattern to a Cooper-pair pairing mechanism (renormalization of the Josephson potential with an added $\cos(2\varphi)$ contribution) owing to stray inductance in the tracks of the SQUID [47, 40] (inductive energy $\epsilon_{L}\gg E_{J}$, see Fig. 2a). Including this mechanism in our model, we quantitatively reproduce all data presented in Fig. 2. The fitted parameters, presented in Table 1, are reasonably close to design values.
@@ III.1 Preparing a Fock state and measuring its occupation
In detail, the states are prepared with pulse sequences combining $\pi$-pulses applied on the single-photon transitions $|0\rangle_{a}\leftrightarrow|1\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|2\rangle_{a}$ (direct charge drive) and/or on the three-photon transitions $|0\rangle_{a}\leftrightarrow|3\rangle_{a}$ and $|1\rangle_{a}\leftrightarrow|4\rangle_{a}$ (flux drive of the ATS). The duration of these pulses being on par with the coherence time of the transitions, the preparation is imperfect. It is nonetheless sufficient to estimate the lifetime of each state, as detailed in the next section. As for the readout, the ground state can be directly distinguished from excited states through the pull exerted by excited states on the readout mode resonance frequency. However, their respective pulls saturate rapidly with increasing state index, making them hard to distinguish from one another. We circumvent this shortcoming with the scheme depicted in Fig. 3a. To estimate the occupation of state $|n\rangle_{a}$ ($n\geq 1$), we apply or not a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long pulse. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the ground state is delayed compared to the decay from $|n\rangle_{a}$, we detect a time-shifted signal in the reflected field $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse with respect to the field $s_{n}^{OFF}(t)$ detected in absence of the pulse [40]. We integrate the detected fields with a time-envelope $\mathcal{E}(t)$ chosen to maximize the signal-to-noise ratio of the differential signal $S_{n}=\int_{t}\Re\Big{(}\big{(}s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big{)}\mathcal{E}(t)\Big{)}\mathrm{d}t$ between these two situations [40]. Since the $\pi$-pulse only modifies the memory state when $|n\rangle_{a}$ is occupied, the integrated signal $S_{n}$ is proportional to the occupation of state $|n\rangle_{a}$ at the instant when the pulse is delivered, even if population transfer is imperfect. Note that both for state preparation and detection, we favor pulses applied to three-photon transitions to ones applied to single-photon transitions as the detuning between neighboring three-photon transitions is larger, allowing for stronger drives.
@@
Before applying these preparation and measurement methods to characterize the memory population dynamics in presence of a pump, we report that some of the memory transitions—driven either during preparation or measurement, involving a single photon or three—were frequently and transiently observed to shift or split [40]. We attribute this effect to hybridization of the memory to a bath of Two-Level-Systems (TLS). Such TLS baths, whose frequency landscape reconfigure on timescales of the order of a second, are frequently observed in superconducting circuits [49, 50]. While the exact coupling mechanism and density of TLS’s was not characterized in our experiment, at least one of the memory transitions was typically affected by a TLS at any given time so that waiting for the bath to reconfigure in a more favorable configuration (either passively or cycling the system to room temperature) was not a viable option. Instead, we use an off-resonant flux drive of large amplitude to shift the frequency of the memory transitions away from any resonant coupling with TLS’s [40]. The frequency of this drive was chosen to minimize decoherence induced on the memory, but still leads to degradation of its energy relaxation and dephasing times 555 Note that in principle, the memory coherence should not be affected by this drive as the ATS remains at a dynamical sweet-spot, in the sense that its susceptibility to magnetic flux drifts averages out over a period of the drive. The datasets presented in the following section were acquired in presence of such a drive.
@@ III.2 Characterization of the four-photon dissipation channel
In order to quantitatively relate the largest fitted values of $g_{4}$ to four-photon dissipation rates, one needs to go beyond the simple model reduction allowed by adiabatic elimination of the buffer. We instead extract the lifetime of state $|4\rangle_{a}$ through spectral analysis of a reduced Liouvillian capturing the system dynamics [40]. We report the extracted decay rate $\Gamma_{4}$ in Fig. 4 (filled red circles). Following an initial quadratic increase of $\Gamma_{4}$ against pump amplitude—also predicted by the adiabatic elimination of the buffer—we observe a saturation of $\Gamma_{4}$ near its theoretical maximum value $\kappa_{b}/2$ set by the buffer damping rate [27]. In this regime, engineered four-photon dissipation dominates the dynamics, inducing decay of population from state $|4\rangle_{a}$ at a rate that exceeds that induced by native single-photon dissipation by a factor 10. As a sanity check, we also record, under the same conditions, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ and report the corresponding decay rates $\Gamma_{n}$ in Fig. 4. While these should nominally not be affected by the pump (see Fig. 1b), we observe a small but clear increase with pump amplitude. Such behavior was reported in similar parametrically pumped systems [23, 25], and and may stem from spurious near-resonant parametric processes [52, 53, 54, 55]. In our case, we note that a two-to-one photon exchange process between the memory and the buffer, resonant at $2\omega_{b}-\omega_{a}$, is detuned by approximately 200 MHz from the target process. We estimate that this off-resonantly driven process results in an effective pump-dependent single-photon dissipation channel for the memory, contributing marginally to the observed degradation of the low-energy levels lifetime (dissipation rate in the $0-10$ kHz range for the considered pump amplitudes). In any case, the spurious increase of $\Gamma_{n}$ for $n<4$ remains an order of magnitude lower than the recorded increase of $\Gamma_{4}$.
IV Discussions
IV.1 High-order dissipators from high-order interactions
In earlier experimental works realizing high-order dissipation channels, these were obtained from low-order interactions. Among other examples, autonomous parity recovery in a superconducting resonator was achieved by combining dispersive interactions and a set of single and two-photon drives [37], autonomous stabilization of the GKP code was achieved by combining Rabi interactions and sequences of single-photon drives [17, 19, 56], and a four-to-two photon exchange interaction—which could be used to engineer a four-photon dissipation channel—was obtained by cascading off-resonant two-to-one photon exchange interactions in a Raman configuration [57] [^6] In that work, the auxiliary mode was not designed to dissipate excitations at a sufficiently large rate to turn the interaction into an effective four-photon dissipation channel.. We argue that this approach can lead to the opening of spurious dissipation channels entailed by each low-order primitive involved in the generation of effective high-order interactions. For instance, in the Raman scheme, finite detuning of the pumps to two-to-one photon transitions can excite levels which should nominally be only virtually excited. Such excitations would directly lead to a logical flip of the encoded cat-qubit if the engineered dissipation channel were employed toward this goal.
In principle, low-order interactions can still be safely used leveraging a biased-noise auxiliary mode [59, 60, 61], error-transparent interactions [62, 63, 64, 65, 29] or carefully controlling the noise spectrum of the dissipative environment [66], but the experimental implementation of these ideas is a major challenge. On the other hand, in our approach, the desired high-order dissipation is directly obtained from high-order interactions, acting on the memory mode via operators which commute with the logical operators of the encoded qubit (the interaction Hamiltonian ${{\bf a}}^{4}{\bf b}^{\dagger}+\mathrm{h.c.}$ commutes with the four-legged cat code stabilizers and its Pauli operators). Therefore, even in presence of noise affecting the buffer, evolutions induced by these interactions do not affect the encoded information. Moreover, it is straightforward to adapt the circuit and method presented in this work to activate even higher-order dissipators.
IV.2 Parameter regime
As discussed in this paper introduction, the memory and buffer phase fluctuations across the ATS should be large in order for the rate of the activated four-to-one photon exchange to be significant. However, the memory phase fluctuations cannot be chosen arbitrarily large when stabilizing four-legged cats [40]. Indeed, for a pump $\epsilon(t)=\xi e^{i(4\omega_{a}-\omega_{b})t}+\mathrm{c.c}.$ resonant with the four-to-one photon transition, the RWA yields corrections to the exchange interaction stemming from high-order terms in the ATS sine Hamiltonian (see Eq. (2)):
In this expression, :_::\_:: _ : denotes the normal ordering of an operator, $g_{4}=2\frac{E_{J}}{\hbar 4!}\xi e^{-(\varphi^{{ZPF}^{2}}_{a}+\varphi^{{ZPF}^{2}}_{b})/2}\varphi^{{ZPF}^{4}}_{a}\varphi^{ZPF}_{b}$ and $P_{a}$ and $P_{b}$ are polynomials boiling down to unit factors if the SQUID potential is truncated to leading order. These corrections yield a state-dependent four-to-one photon exchange rate which can result in instabilities when stabilizing a cat with size $\alpha\gtrsim 1/\varphi_{a}^{ZPF}$ [40, 21]. We note that this effect also arises for two-legged cats, and is one of the limiting factors to their stabilization rate  [^7] Large dissipation rates are desired for fast adiabatic control of the cat state [75] but are not needed to achieve macroscopic bit-flip times [26].. On the upside, it was recently proposed to leverage the same effect to increase the number of attractors in phase-space—thus stabilizing higher-dimension cat codes—or to stabilize more complex rotation-symmetric codes with advantageous error-correcting capabilities [21].
In circuit design, we target the value of $\varphi_{a}^{ZPF}=0.45$, for which numerical simulations showed that cat states containing more than 10 photons could be stabilized with no visible impact from high-order corrections to the four-to-one photon interaction. Note that owing to the exponential scaling of the cat bit-flip time against the cat size, this mean photon number is typically sufficient for cats to reach macroscopic bit-flip times [27]. For the buffer, we target $\varphi_{b}^{ZPF}=0.55$ (significantly larger than the measured value). This is lower than the unit value maximizing the four-to-one photon exchange rate $g_{4}$. However, reaching this regime would have required to increase significantly the mode impedance for only a modest increase of the exchange rate. As for the readout mode, its phase fluctuations across the ATS are an order of magnitude lower owing to a weaker participation of the ATS. This mode was originally intended to provide a photon number resolving measurement by activating a term of the form ${\bf a}^{\dagger}{\bf a}({\bf r}+{\bf r}^{\dagger})$ (sometimes referred to as "longitudinal" coupling [68]) by pumping the ATS at the readout mode frequency [27]. In this context, setting $\varphi_{r}^{ZPF}\ll\varphi_{a}^{ZPF}$ ensures that unwanted cubic interactions of readout photons, which are also activated by the pump, are significantly weaker than the desired term. In practice, the memory and readout modes were found to be dispersively coupled due to circuit non-idealities such that dispersive measurement was sufficient to characterize the state of the memory (see Sec. III.1).
IV.3 Spurious non-linearities and mitigation measures
In the introduction of this paper, we discussed how an idealized circuit comprising a SQUID placed in series with two resonators enabled dissipation engineering while preserving the target mode linearity. Various non-idealities can break this property and induce anharmonicity. We here outline how the observed memory anharmonicity can be traced back to stray inductance in the SQUID and give possible mitigation measures for future experiments. We refer the reader to the paper Supplemental Materials for a detailed analysis of the various non-linear mechanisms arising in our circuit.
The memory anharmonicity pattern presented in Fig. 2d significantly differs from that of a Kerr oscillator, for which a constant anharmonicity is expected. It is characteristic of a strongly non-linear, i.e. high-order, perturbation of weak amplitude. Such a perturbation impacts low-energy states but is overcome by the quadratic part of the Hamiltonian at high energy. This disqualifies the chain of Josephson junctions forming the circuit superinductors—expected to induce a negative quartic non-linearity—as the dominant source of non-linearity. Imbalance of the SQUID junctions—often responsible for spurious nonlinearities in circuits comprising an ATS [25]—is here negligible and cannot explain the magnitude of the anharmonicity. On the other hand, the anharmonicity pattern is quantitatively reproduced when fitted with a small second harmonic contribution to the potential of each Josephson junction forming the SQUID. Including this contribution, the ATS Hamiltonian (2) becomes
${\bf H}_{J}(t)$ $=-E_{J}\cos\big{(}\bm{\varphi}-\frac{\Phi_{1}}{\varphi_{0}}\big{)}-E_{J}\cos\big{(}\bm{\varphi}+\frac{\Phi_{2}}{\varphi_{0}}\big{)}$ (5)
$\qquad-E_{J_{2}}\cos\big{(}2(\bm{\varphi}-\frac{\Phi_{1}}{\varphi_{0}})\big{)}-E_{J_{2}}\cos\big{(}2(\bm{\varphi}+\frac{\Phi_{2}}{\varphi_{0}})\big{)}$
$=2E_{J}~{}\sin\big{(}\epsilon(t)\big{)}\sin(\bm{\varphi})$
$\qquad-2E_{J_{2}}~{}\cos\big{(}2\epsilon(t)\big{)}\cos(2\bm{\varphi})$
We immediately see that the added contribution does not cancel at the ATS working point, even in absence of pump ($\epsilon=0$). Moreover, the vacuum phase fluctuations of each mode are effectively doubled and approach unity for this term, such that the cosine cannot be approximated by a low-order polynomial. For the fitted value $E_{J_{2}}=E_{J}/500$, we thus find a strongly non-linear perturbation of weak amplitude as sought.
A known mechanism renormalizing the Josephson energy with a second harmonic stems from small stray inductances in the arms of the SQUID [47, 54]. For series inductance with inductive energy $\epsilon_{L}\gg E_{J}$ (see Fig. 2a), the amplitude of the second harmonic is $E_{J_{2}}=-E_{J}^{2}/(4\epsilon_{L})$ [40]. In our case, the observed amplitude corresponds to an inductance value of $\varphi_{0}^{2}/\epsilon_{L}=85$ pH matching our estimate for the geometric inductance of the tracks of the SQUID. Note that an alternative mechanism yielding harmonics in the Josephson potential—stemming from pinholes in the Josephson junctions oxide barrier—was recently proposed to explain reported anomalies in transmon energy level structures [69]. The amplitude of the second harmonic found in our experiment is smaller than the values reported in that work and can be entirely attributed to stray circuit inductances. This different behavior may be explained by a thicker oxide barrier in our junctions [40].
In a future iteration of the experiment, stray inductances could be significantly reduced with simple design changes. Indeed, the rather large SQUID loop in the present experiment (dimensions $\sim 50\times 50~{}\mathrm{\mu m}^{2}$) was chosen to accommodate at its center the long chain of junctions forming the shunt inductor of the ATS. However, physically placing this chain in parallel but outside of the SQUID loop would yield an equivalent circuit—simply requiring to reallocate bias fluxes between loops—while allowing to reduce the SQUID perimeter down to $\sim 10~{}\mu$m as routinely achieved in flux qubits [70]. Combined with a reduction of the junctions energy to $E_{J}=h\times 5~{}$GHz—as was originally intended in design—we estimate that the anharmonicity of the memory mode will become sufficiently low to be overcome by the four-photon dissipation.
Finally, we note that while our experiment is particularly sensitive to non-linearity resulting from various circuit non-idealities (imbalance of the SQUID, non-linearity of the superinductors, harmonics in the Josephson potential), these mechanisms also exist in experiments aiming at dissipatively stabilizing other bosonic codes with similar circuits. While these experiments typically operate in a lower impedance regime in order to to suppress the anharmonicity inherited by the memory (dominant non-linear term scaling as ${\varphi_{a}^{ZPF}}^{4}\propto Z_{a}^{2}$), the rate of desired interactions decreases in the same limit. Therefore, even when targeting lower-order dissipation channels scaling more favorably with the modes impedance, the optimal parameter choice results from a tradeoff between the rate of engineered dissipation and the strength of spurious non-linearities. Our platform is an ideal testbed to investigate this tradeoff. Our experiment also highlights that the anharmonicity inherited from different mechanisms cannot be boiled down to the amplitude of a single Kerr-like term. As a result, compensating the effect of one mechanism with another as was recently realized [54] can only lead to imperfect cancellation. An intriguing exception detailed in this paper Supplemental materials is the possibility to perfectly cancel, in the RWA, the anharmonicity induced by a weak but non-zero imbalance of the SQUID with a dynamical mechanism involving off-resonant pumps. In future experiments, this scheme should prove instrumental in order not to rely on perfect SQUID symmetry as was achieved in the present experiment.
V Conclusion and outlook